OpenStudy (anonymous):

Find a rational number and an irrational number that are between 5.2 and 5.5. Include the decimal approximation of the irrational number to the nearest hundredth.

OpenStudy (anonymous):

@juliannescriv

OpenStudy (anonymous):

i'll medal n fan you if someone helps please.

OpenStudy (zarkon):

what have you tried?

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@Donblue22 @johnweldon1993

OpenStudy (anonymous):

i cant do it ive tried its a pretest so idk this stuff

OpenStudy (anonymous):

pls help me somebody

OpenStudy (zarkon):

do you know what rational and irrational numbers are?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

then I don't understand what you are having issues with then

OpenStudy (anonymous):

rational can be written as a fraction and irrational are endless numbers

OpenStudy (anonymous):

i dont understand how to find it....

OpenStudy (anonymous):

just bc i know what the terms are doesnt necessarily mean i understand the question

OpenStudy (zarkon):

rational numbers can also be written as a terminating or repeating decimal

OpenStudy (anonymous):

okay

OpenStudy (zarkon):

5.2 is rational....5.5 is also a rational number

OpenStudy (anonymous):

okay. so 5.3 would be rational too?

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

5.3=53/10

OpenStudy (anonymous):

okay. how about an irrational number? i get confused on that part.

OpenStudy (zarkon):

it would be easy to construct an irrational number but I doubt that is what they want since they only want 2 decimal digits.

OpenStudy (zarkon):

so use the number \(\pi\) find a rational number that when you multiply it by \(\pi\) you get a number between 5.2 and 5.5

OpenStudy (anonymous):

thank u!!